# TaP: A Taxonomy-Guided Framework for Automated and Scalable Preference Data Generation

### v3

@@ Abstract
Conducting supervised fine-tuning and preference fine-tuning on large language models (LLMs) requires high-quality datasets to improve their ability to follow instructions and align with human preferences and values. However, constructing such datasets is resource-intensive, and most available datasets for supervised and preference fine-tuning are in English. To address these challenges, we propose the Taxonomy-Guided Preference Data Generation (TaP) framework, which facilitates automated and scalable construction of preference datasets across various languages. TaP is grounded in a structured taxonomy that allows fine-grained control over dataset composition, thereby ensuring both diversity and comprehensive coverage. We employ TaP-generated datasets to perform supervised and preference fine-tuning on various LLMs. Experimental results demonstrate that LLMs trained on TaP-generated datasets outperform those trained on existing open-source datasets. Remarkably, LLMs trained on TaP-generated datasets surpass the performance of those trained on an open-source dataset that is 180 times larger.
@@ 1 Introduction
Large language models (LLMs) typically undergo three primary training stages: (1) pre-training, during which LLMs are trained on extensive and diverse datasets encompassing multiple languages and modalities (DBLP:journals/corr/abs-2303-18223; DBLP:journals/corr/abs-2412-15115; DBLP:journals/corr/abs-2407-21783; DBLP:journals/corr/abs-2412-19437); (2) supervised fine-tuning, in which LLMs are fine-tuned using prompt-response pairs authored by humans or LLMs (DBLP:journals/corr/abs-2308-10792); and (3) preference fine-tuning, wherein preference data annotated by humans or LLMs are employed to conduct RLHF or alternative methods such as DPO and KTO (DBLP:conf/nips/Ouyang0JAWMZASR22; DBLP:conf/nips/RafailovSMMEF23; DBLP:journals/corr/abs-2402-01306). The latter two stages are essential for improving LLMs’ ability to follow instructions and generate responses that are aligned with human preferences and values.
@@
However, datasets for supervised and preference fine-tuning usually require human annotation, a process that is time-consuming and expensive. These constraints significantly hinder the scalability of dataset expansion, thereby limiting the volume of available training data. A common alternative involves leveraging LLMs to generate synthetic data, reducing dependence on human labor and enhancing scalability. Nevertheless, naive prompting often results in homogeneous data with limited coverage and pronounced task or domain imbalances, wherein certain domains or tasks are overrepresented while others remain underrepresented (DBLP:conf/acl/WangKMLSKH23; DBLP:journals/corr/abs-2306-11644; DBLP:conf/iclr/XuJNDP0L25). Such imbalances can lead LLMs to overfit to the overrepresented domains or tasks, thereby weakening the models’ generalizability in real-world applications. Consequently, ensuring both diversity and comprehensive coverage in synthetic data generation remains a critical challenge.
@@
Furthermore, compared with open-source datasets for supervised fine-tuning, there are significantly fewer datasets for preference fine-tuning (DBLP:conf/icml/CuiY0YH0NXXL0024; DBLP:conf/naacl/WangDZASEDSKSK24; DBLP:journals/corr/abs-2406-08673). Additionally, open-source datasets for both supervised and preference fine-tuning are predominantly available in English. This language bias limits the global applicability of LLMs and marginalizes non-English-speaking communities, preventing them from fully benefiting from these models. While translating English datasets into other languages presents a potential solution, it can introduce translation errors and result in “translationese”, thereby compromising dataset quality (DBLP:conf/acl/RileyCFG20). Given that there are over 7,000 languages worldwide, it is imperative to develop novel approaches and release open-source datasets in diverse languages to better support the latter two training stages of LLMs.
@@
To address these challenges, we propose the Taxonomy-Guided Preference Data Generation (TaP) framework, which enables automated and scalable generation of preference data across various languages. The TaP framework leverages a predefined taxonomy to guide LLMs in producing preference data spanning a wide range of domains and tasks. By generating data in accordance with the predefined taxonomy, TaP ensures fine-grained control over dataset composition and facilitates the construction of diverse datasets with comprehensive coverage of the categories defined within the taxonomy. The TaP framework comprises four key steps: (1) designing a taxonomy aligned with the intended application of LLMs—for instance, a taxonomy for general-purpose LLMs can be organized around common use cases such as brainstorming, open-ended question answering, and summarization; (2) prompting LLMs to generate prompts corresponding to the categories specified in the taxonomy; (3) employing multiple LLMs to generate diverse responses for each prompt; and (4) utilizing LLMs to evaluate and rank these responses according to predefined evaluation criteria. A diagram of the preference dataset construction process using the TaP framework is presented in Figure 1.
@@
Motivated by the fundamental role of knowledge in decision-making, problem-solving, and adaptation to new environments, as well as by empirical findings indicating that LLMs benefit from diverse and knowledge-rich training data (DBLP:journals/corr/abs-2309-05463; lozhkov2024fineweb-edu), we propose leveraging a comprehensive taxonomy that encompasses a broad spectrum of knowledge and skills essential to human expertise to guide the generation of such preference data using LLMs. Through careful manual analysis, we identify undergraduate program catalogs as a valuable resource for constructing such a taxonomy. These catalogs serve as structured and comprehensive repositories of essential knowledge required for human learning. Considering that undergraduate education constitutes a foundational component of higher education systems worldwide, and is available in most countries, [^1] https://en.wikipedia.org/wiki/Undergraduate_education employing undergraduate program catalogs enhances the adaptability of the TaP framework across various languages. For cases where an undergraduate program catalog is not available in a given language, we discuss alternative approaches for constructing the taxonomy in Appendix A.
@@
Given the abundance of publicly available preference datasets in English and the relative scarcity of such datasets in other languages, we apply the TaP framework to generate preference data in Chinese. Specifically, we adopt the Undergraduate Program Catalog of Regular Higher Education Institutions of China [^2] http://www.moe.gov.cn/srcsite/A08/moe_1034/s4930/202403/W020240319305498791768.pdf as the guiding taxonomy for preference data generation. Manual inspection of the generated data reveals several desirable properties: (1) rich in knowledge, (2) high information density, and (3) broad coverage across domains and tasks. An illustrative example of a prompt and corresponding response is presented in Figure 3 of Appendix B to highlight these properties. We conduct both supervised fine-tuning and preference fine-tuning, including PPO and DPO, on LLMs using the TaP-generated data. Experimental results demonstrate that LLMs fine-tuned with TaP-generated data outperform those trained on existing open-source datasets. Our contributions can be summarized as follows:
@@
  • We propose the TaP framework, which enables automated, scalable preference data generation across various languages and offers fine-grained control over data composition.
@@
  • We apply TaP to generate high-quality preference data in Chinese, which can be used for both supervised and preference fine-tuning (e.g., PPO, DPO, GRPO). Additionally, the annotations produced by LLMs during the data generation process can be used to fine-tune open-source LLMs as judges (DBLP:conf/iclr/WangYYZYW0J000024; DBLP:conf/emnlp/KimSLLSWNL0S24; DBLP:conf/iclr/ZhuWW25), thereby reducing reliance on costly proprietary APIs and the risk of data leakage.
@@
  • We conduct extensive experiments using TaP-generated preference data to train five open-source LLMs ranging in size from 3B to 14B parameters, across three model families: LLaMA-3.1, Qwen2.5, and Gemma-2. The experimental results demonstrate that models trained on TaP-generated data outperform those trained on existing open-source datasets.
@@ 2 Related Work
Our research is closely related to studies on synthetic data generation using LLMs. Pioneering studies in this field have leveraged the strong in-context learning capabilities of LLMs by providing several demonstrations to guide the generation of synthetic samples (DBLP:conf/acl/HonovichSLS23; DBLP:conf/acl/WangKMLSKH23), typically requiring a small set of seed samples. Following this line of research, the follow up research usually necessitating a small number of seed samples to guide the LLMs in producing synthetic samples (DBLP:conf/iclr/XuSZG0FTLJ24; DBLP:journals/corr/abs-2306-02707; DBLP:conf/iclr/LiYZSLZWL24; DBLP:conf/nips/YueZZC24).
In contrast, TaP employs a taxonomy to guide LLMs in autonomously generating synthetic data, eliminating the need for seed samples authored by humans or LLMs. While TaP shares the goal of constructing preference datasets with UltraFeedback (DBLP:conf/icml/CuiY0YH0NXXL0024) and the HelpSteer series (DBLP:conf/naacl/WangDZASEDSKSK24; DBLP:journals/corr/abs-2406-08673), it differs in two key aspects: (1) TaP can produce prompts from scratch without relying on existing instruction datasets. This enables fine-grained control over dataset composition and enhances its applicability across various languages. Conversely, DBLP:conf/icml/CuiY0YH0NXXL0024 and DBLP:conf/naacl/WangDZASEDSKSK24; DBLP:journals/corr/abs-2406-08673 construct preference datasets by collecting prompts from existing instruction datasets. (2) TaP is applied to construct preference datasets in Chinese, addressing the scarcity of such data in that language. In comparison, DBLP:conf/icml/CuiY0YH0NXXL0024 and DBLP:conf/naacl/WangDZASEDSKSK24; DBLP:journals/corr/abs-2406-08673 primarily focus on constructing preference datasets in English.
@@ 3 Preference Dataset Curation
The TaP framework for automated and scalable preference data generation comprises four key steps: (1) Taxonomy Specification, wherein a structured taxonomy is constructed to comprehensively cover the common use cases of LLMs; (2) Prompt Generation, wherein LLMs are guided to produce prompts corresponding to the categories defined in the taxonomy; (3) Response Generation and Selection, wherein various LLMs generate responses to each prompt, and a subset is selected to balance annotation cost and response diversity; and (4) Fine-grained Preference Annotation, wherein LLMs evaluate and score the selected responses across multiple dimensions. Figure 1 illustrates the entire process of constructing the preference dataset using the TaP framework.
@@ 3.1 Taxonomy Specification
We adopt the Undergraduate Program Catalog of Regular Higher Education Institutions of China as the taxonomy to guide LLMs in generating preference data in Chinese. This catalog comprehensively lists all undergraduate subjects offered by Chinese universities, encompassing a diverse range of academic disciplines, spanning humanities, social sciences, and STEM fields. It comprises 816 distinct subjects, systematically grouped into 93 discipline categories. These categories are further organized into 12 primary degree-conferring disciplines, ensuring a structured organization of undergraduate knowledge.
To construct preference datasets that cover diverse domains and tasks, we have incorporated all subjects specified in the catalog into our taxonomy, with the exception of those in the Foreign Languages and Literatures discipline, which includes numerous low-resource languages. Instead, we have selectively included 12 subjects: the top nine languages, along with Linguistics, Translation, and Business English. Consequently, our taxonomy encompasses a total of 724 subjects, consisting of 12 from the Foreign Languages and Literatures discipline and 712 from other disciplines.
@@ 3.2 Prompt Generation
We adopt a multi-stage prompting strategy, rather than instructing LLMs to generate prompts directly, to enhance both the diversity and coverage of the resulting outputs. In the first stage, LLMs are prompted to generate diverse question types and brief descriptions for each subject within a predefined taxonomy through a three-round interaction. Each round produces a distinct set of question types and descriptions, thereby increasing both the quantity and diversity of the generated outputs. The prompt used in this stage is provided in Figure 4 of Appendix C.1. In the second stage, LLMs refine the question type descriptions from the first stage to improve their clarity and real-world relevance. This refinement aims to enhance the overall quality of the final prompts. The prompt used in this stage is shown in Figure 5 of Appendix C.1. In the third stage, we employ a two-round interaction to address the issue of LLMs occasionally producing incomplete prompts lacking essential contextual information. [^3] For instance, LLMs may generate prompts that request a summary of an article without providing the article itself. In the first round, LLMs generate a candidate prompt. In the second round, they assess whether the prompt contains all necessary information. If a prompt is identified as incomplete, the LLMs regenerate it to ensure completeness. The prompt used for this interaction is illustrated in Figure 6 of Appendix C.1. Additionally, we find that LLMs sometimes generate prompts beyond their capabilities, such as those requiring real-world actions. To mitigate this issue, we employ the LLM as a checker to evaluate whether a given prompt is feasible within its operational constraints. If a prompt is deemed infeasible, it is revised, with up to three regeneration attempts. The evaluation prompt for this step is presented in Figure 7 of Appendix C.1. To verify the reliability of LLMs in identifying such infeasible prompts, we manually reviewed a subset of evaluations generated by GPT-4. The results demonstrate that GPT-4 reliably identifies prompts that exceed its capabilities. Further details are provided in Appendix C.2.
@@
Huozi-RLHF (huozi) 16,918 3,725 69.19 $\pm$ 107.38 216.02 $\pm$ 165.44 Ranking Human Chinese
@@
Table 1: Comparison of open-source preference datasets in Chinese with the dataset constructed using TaP. “TaP (GPT-4)” indicates that the prompts in the dataset were generated by GPT-4. The column labeled “#Prompts” represents the number of unique prompts in each dataset. “Prompt Length” and “Response Length” are expressed in the format $XX\pm YY$, where XXXX denotes the mean number of tokens and YYYY represents the standard deviation. Tokenization was performed using the cl100k_base encoding provided by the tiktoken library. [^5] https://github.com/openai/tiktoken
@@ 3.3 Response Generation and Selection
We select 65 widely used LLMs, ranging in size from 0.5B to 236B parameters, to generate responses for each prompt, thereby enhancing the diversity and coverage of LLM-generated responses. The models span both dense and MoE architectures and include both proprietary and open-source models. The rationale for model selection, along with detailed specifications of the chosen LLMs, is provided in Appendix C.3. Each model is prompted using the system prompt shown in Figure 8 of Appendix C.1, resulting in 65 responses per prompt.
Given that each prompt produces 65 responses, leading to 2,080 pairwise preference pairs to be annotated per prompt. To manage dataset size and reduce annotation costs while preserving response diversity and coverage, we apply K-means clustering to group the responses for each prompt into five clusters. From each cluster, we select a representative response. Specifically, we choose the response closest to the cluster centroid, except in the case of the cluster containing the response generated by GPT-4. In this case, we instead select the GPT-4 response to ensure that at least one high-quality response is included for each prompt. For K-means clustering, we obtain response embeddings by applying mean pooling over the hidden states of all tokens from the final layer of Qwen2-7B.

### v4

@@ Abstract
[…]
@@ 1 Introduction
Large language models (LLMs) typically undergo three primary training stages: (1) pre-training, during which LLMs are trained on extensive and diverse datasets encompassing multiple languages and modalities (Zhao et al., 2023; Yang et al., 2024; Dubey et al., 2024; DeepSeek-AI et al., 2024); (2) supervised fine-tuning, in which LLMs are fine-tuned using prompt–response pairs authored by humans or LLMs (Zhang et al., 2023b); and (3) preference fine-tuning, wherein preference data annotated by humans or LLMs are used to conduct RLHF or alternative methods such as DPO and KTO (Ouyang et al., 2022; Rafailov et al., 2023; Ethayarajh et al., 2024). The latter two stages are essential for improving LLMs’ ability to follow instructions and generate responses aligned with human preferences and values.
@@
However, datasets for supervised and preference fine-tuning often require human annotation, which is time-consuming and expensive. These constraints hinder the scalability of dataset expansion, limiting the amount of available training data. A common alternative is to leverage LLMs to generate synthetic data, thereby reducing reliance on human labor and improving scalability. Nevertheless, naive prompting often yields homogeneous data with limited coverage and pronounced task or domain imbalance, where certain domains or tasks are overrepresented while others remain underrepresented (Wang et al., 2023; Gunasekar et al., 2023; Xu et al., 2025). Such imbalance can cause LLMs to overfit overrepresented domains or tasks, weakening their generalizability in real-world applications. Consequently, ensuring both diversity and comprehensive coverage in synthetic data generation remains a critical challenge.
@@
Furthermore, compared with open-source datasets for supervised fine-tuning, far fewer datasets are available for preference fine-tuning (Cui et al., 2024; Wang et al., 2024d, c). Additionally, open-source datasets for both supervised and preference fine-tuning are predominantly in English. This language bias limits the global applicability of LLMs and marginalizes non-English-speaking communities, preventing them from fully benefiting from these models. Although translating English datasets into other languages is a potential solution, it can introduce translation errors and produce “translationese”, thereby compromising dataset quality (Riley et al., 2020). Given that there are over 7,000 languages worldwide, it is imperative to develop novel approaches and release open-source datasets in diverse languages to better support the latter two training stages of LLMs.
@@
To address these challenges, we propose the Taxonomy-Guided Preference Data Generation (TaP) framework for automated, scalable preference data generation across languages. TaP leverages a predefined taxonomy to guide LLMs in producing preference data spanning diverse domains and tasks. By generating data according to the taxonomy, TaP provides fine-grained control over dataset composition and facilitates the construction of diverse datasets with broad coverage of the taxonomy categories. The TaP framework comprises four steps: (1) designing a taxonomy aligned with the target application of LLMs (e.g., for general-purpose LLMs, organizing the taxonomy around use cases such as brainstorming, open-ended question answering, and summarization); (2) prompting LLMs to generate prompts for the categories specified in the taxonomy; (3) employing multiple LLMs to generate diverse responses for each prompt; and (4) using LLMs to evaluate and rank responses according to predefined criteria. A diagram of the dataset construction process using the TaP framework is shown in Figure 1.
@@
Motivated by the fundamental role of knowledge in decision-making, problem-solving, and adaptation to new environments, as well as by empirical findings indicating that LLMs benefit from diverse and knowledge-rich training data (Li et al., 2023; Lozhkov et al., 2024), we propose leveraging a comprehensive taxonomy, which encompasses a broad spectrum of knowledge and skills essential to human expertise, to guide the generation of such preference data using LLMs. Through careful manual analysis, we identify undergraduate program catalogs as a valuable resource for constructing such a taxonomy, as they are structured, comprehensive repositories of essential knowledge for human learning. Considering that undergraduate education constitutes a foundational component of higher education systems worldwide and is available in most countries, [^1] https://en.wikipedia.org/wiki/Undergraduate_education employing undergraduate program catalogs enhances the adaptability of the TaP framework across various languages. When such catalogs are unavailable in a given language, we discuss alternative approaches for constructing the taxonomy in Appendix A.
@@
Given the abundance of publicly available preference datasets in English and the relative scarcity of such datasets in other languages, we apply the TaP framework to generate preference data in Chinese. Specifically, we adopt the Undergraduate Program Catalog of Regular Higher Education Institutions of China [^2] http://www.moe.gov.cn/srcsite/A08/moe_1034/s4930/202403/W020240319305498791768.pdf as the guiding taxonomy for preference data generation. Manual inspection of the generated data reveals several desirable properties: (1) rich in knowledge, (2) high information density, and (3) broad coverage across domains and tasks. An illustrative example of a prompt and corresponding response is presented in Figure 3 in Appendix B to highlight these properties. We conduct both supervised and preference fine-tuning, including PPO and DPO, on LLMs using the TaP-generated data. Experimental results demonstrate that LLMs fine-tuned with TaP-generated data outperform those trained on existing open-source datasets. Our contributions can be summarized as follows:
@@
  • We propose the TaP framework, which enables automated, scalable preference data generation across languages and provides fine-grained control over dataset composition.
@@
  • We apply TaP to generate high-quality Chinese preference data for supervised and preference fine-tuning (e.g., PPO, DPO, GRPO). Additionally, the annotations produced by LLMs during data generation can be used to fine-tune open-source LLMs as judges (Wang et al., 2024b; Kim et al., 2024; Zhu et al., 2025b), reducing reliance on costly proprietary APIs and the risk of data leakage.
@@
  • We conduct extensive experiments using TaP-generated preference data to train five open-source LLMs (3B–14B) from three model families: LLaMA-3.1, Qwen2.5, and Gemma-2. The results show that models trained on TaP-generated data outperform those trained on existing open-source datasets.
@@ 2 Related Work
Our research is closely related to studies on synthetic data generation using LLMs. Pioneering studies in this field have leveraged the in-context learning capabilities of LLMs by providing demonstrations to guide the generation of synthetic samples (Honovich et al., 2023; Wang et al., 2023), typically requiring a small set of seed samples. Following this line of research, subsequent work has typically required a small number of seed samples to guide LLMs in producing synthetic samples (Xu et al., 2024; Mukherjee et al., 2023; Li et al., 2024; Yue et al., 2024; Zhu et al., 2025a; Ren et al., 2025).
In contrast, TaP uses a taxonomy to guide LLMs in autonomously generating synthetic data, eliminating the need for seed samples authored by humans or LLMs. While TaP shares the goal of constructing preference datasets with UltraFeedback (Cui et al., 2024) and the HelpSteer series (Wang et al., 2024d, c), it differs in two key aspects. First, TaP can produce prompts from scratch without relying on existing instruction datasets, enabling fine-grained control over dataset composition and improving its applicability across languages. By contrast, Cui et al. (2024) and Wang et al. (2024d, c) construct preference datasets by collecting prompts from existing instruction datasets. Second, TaP constructs Chinese preference datasets to address the scarcity of such data, whereas Cui et al. (2024) and Wang et al. (2024d, c) mainly focus on English.
@@ 3 Preference Dataset Curation
The TaP framework for automated and scalable preference data generation comprises four key steps: (1) Taxonomy Specification, which constructs a structured taxonomy to comprehensively cover common use cases of LLMs; (2) Prompt Generation, which guides LLMs to produce prompts corresponding to the categories defined in the taxonomy; (3) Response Generation and Selection, in which multiple LLMs generate responses to each prompt and a subset is selected to balance annotation cost and response diversity; and (4) Fine-grained Preference Annotation, in which LLMs evaluate and score the selected responses across multiple dimensions. Figure 1 illustrates the entire process of constructing the preference dataset using the TaP framework.
@@ 3.1 Taxonomy Specification
We adopt the Undergraduate Program Catalog of Regular Higher Education Institutions of China as the taxonomy for guiding LLMs to generate Chinese preference data. This catalog lists all undergraduate subjects offered by Chinese universities, spanning a wide range of disciplines across the humanities, social sciences, and STEM fields. It comprises 816 distinct subjects grouped into 93 discipline categories. These categories are further organized into 12 primary degree-conferring disciplines, providing a structured organization of undergraduate knowledge.
To construct preference datasets covering diverse domains and tasks, we incorporate nearly all subjects specified in the catalog into our taxonomy, except those in the Foreign Languages and Literatures discipline, which includes many low-resource languages. Instead, we selectively include 12 subjects: the top nine languages, along with Linguistics, Translation, and Business English. Consequently, our taxonomy contains 724 subjects in total: 12 from Foreign Languages and Literatures and 712 from other disciplines.
@@ 3.2 Prompt Generation
We adopt a multi-stage prompting strategy, rather than instructing LLMs to generate prompts directly, to improve both the diversity and coverage of the resulting outputs. In the first stage, LLMs are prompted to generate diverse question types and brief descriptions for each subject within a predefined taxonomy through a three-round interaction. Each round produces a distinct set of question types and descriptions, thereby increasing both the quantity and diversity of the generated outputs. The prompt used in this stage is provided in Figure 4 of Appendix C.1. In the second stage, LLMs refine the question-type descriptions from the first stage to improve their clarity and real-world relevance. This refinement aims to improve the quality of the model-generated prompts. The prompt used in this stage is shown in Figure 5 of Appendix C.1. In the third stage, we employ a two-round interaction to address the issue that LLMs occasionally produce incomplete prompts that lack essential contextual information. [^3] For instance, LLMs may generate prompts that request a summary of an article without providing the article itself. In the first round, LLMs generate a candidate prompt based on the question types and refined descriptions. In the second round, they assess whether the prompt contains all necessary information. If a prompt is identified as incomplete, the LLMs regenerate it to ensure completeness. The prompt used for this interaction is illustrated in Figure 6 of Appendix C.1. Additionally, we find that LLMs sometimes generate prompts beyond their capabilities, such as those requiring real-world actions. To mitigate this issue, we employ an LLM as a checker to evaluate whether a given prompt is feasible within its operational constraints. If a prompt is deemed infeasible, it is revised, with up to three regeneration attempts. The evaluation prompt for this step is presented in Figure 7 of Appendix C.1. To verify the reliability of LLMs in identifying such infeasible prompts, we manually reviewed a subset of evaluations generated by GPT-4. The results demonstrate that GPT-4 reliably identifies prompts that exceed its capabilities. Further details are provided in Appendix C.2.
@@
Huozi-RLHF (Huozi-Team, 2024) 16,918 3,725 69.19 $\pm$ 107.38 216.02 $\pm$ 165.44 Ranking Human Chinese
@@
Table 1: Comparison of open-source Chinese preference datasets with the dataset constructed using TaP. “TaP (GPT-4)” indicates that the prompts in the dataset were generated by GPT-4. The column labeled “#Prompts” denotes the number of unique prompts in each dataset. “Prompt Length” and “Response Length” are reported as $XX\pm YY$, where $XX$ is the mean number of tokens and $YY$ is the standard deviation. Tokenization uses the cl100k_base encoding provided by the tiktoken library. [^5] https://github.com/openai/tiktoken
@@ 3.3 Response Generation and Selection
We select 65 widely used LLMs (0.5B–236B parameters) to generate responses for each prompt, increasing response diversity and coverage. The models span dense and MoE architectures and include proprietary and open-source models. The rationale for model selection and detailed model specifications are provided in Appendix C.3. Each model uses the system prompt shown in Figure 8 in Appendix C.1, producing 65 responses per prompt.
Because each prompt yields 65 responses, it produces 2,080 (65 $\times$ 64 / 2) pairwise comparisons for annotation. To reduce annotation cost while maintaining response diversity and coverage, we apply K-means clustering to partition the responses for each prompt into five clusters and select one representative response per cluster. Concretely, we choose the response closest to each cluster centroid, except for the cluster containing the GPT-4 response; in that case, we select the GPT-4 response to ensure that each prompt includes at least one high-quality response. For clustering, we compute response embeddings by mean-pooling token representations from the final layer of Qwen2-7B.